### v1

@@ Abstract
The proliferation of textual data on the Internet presents a unique opportunity for institutions and companies to monitor public opinion about their services and products. Given the rapid generation of such data, the text stream mining setting, which handles sequentially arriving, potentially infinite text streams, is often more suitable than traditional batch learning. While pre-trained language models are commonly employed for their high-quality text vectorization capabilities in streaming contexts, they face challenges adapting to concept drift—the phenomenon where the data distribution changes over time, adversely affecting model performance. Addressing the issue of concept drift, this study explores the efficacy of seven text sampling methods designed to selectively fine-tune language models, thereby mitigating performance degradation. We precisely assess the impact of these methods on fine-tuning the SBERT model using four different loss functions. Our evaluation, focused on Macro F1-score and elapsed time, employs two text stream datasets and an incremental SVM classifier to benchmark performance. Our findings indicate that Softmax loss and Batch All Triplets loss are particularly effective for text stream classification, demonstrating that larger sample sizes generally correlate with improved macro F1-scores. Notably, our proposed WordPieceToken ratio sampling method significantly enhances performance with the identified loss functions, surpassing baseline results.
@@ 1 Introduction
Automatically learning from textual data leveraging machine learning mechanisms brings several challenges in batch processing, such as text standardization and vectorization. Text vectorization plays an important role since most machine learning methods expect numeric vectors as input. Traditional vector representations, such as Bag-of-Words (BOW) [11] and Term Frequency - Inverse of Document Frequency (TF-IDF) [20], can generate very-high-dimensional vectors, which can be disadvantageous to the machine learning model, increasing the computational cost.
@@
The contributions of this paper are 4-fold: (a) an extensive comparison among text sampling methods for fine-tuning purposes; (b) an analysis of the impact of the sampling methods considering the text stream setting; (c) an evaluation of loss functions for fine-tuning SBERT, and (d) a novel textual sampling method based on the ratio between Wordpieces and tokens of a text. The term Wordpieces represents a subword partition system present in BERT [7] that allows handling out-of-vocabulary tokens.
@@ 2 Background
This section presents core concepts for understanding this paper. In particular, we introduce text stream mining, SentenceBERT, and its fine-tuning process.
@@ 2.1 Text Stream Mining
According to Bifet et al. [3], data streams are “an abstraction that allows real-time analytics”. In a data stream, the items arrive individually or in small sequential batches, and the stream itself can be infinite [3, 8]. To learn from data streams, different learning approaches have been developed, including, for instance [8]: the ability to learn incrementally, single-pass operations, elimination of input data as soon as possible after learning from it, and consumption of modest resources, i.e., processing power and time.
@@
Frequently, text-related approaches leverage pre-trained language models [9, 24]. Using pre-trained models can help save time since training a language model from scratch is computationally costly in time and resources [22]. In addition, the language model can easily be reused in different scenarios. However, an important drawback is that texts generally suffer from concept drift. Concept drift is a phenomenon frequently observed in real-world datasets and corresponds to changes in data distribution over time [8, 9]. Leveraging pre-trained language models without accounting for concept drift can lead to a decrease in the performance in the downstream task since the texts would be represented using relatively old representations [9].
@@ 2.2.1 Fine-tuning, data preparation, and loss functions
  • Softmax loss (SL) receives text pairs and a label indicating their relationship. Reimers and Gurevych [19] leveraged this loss function for natural language inference, and therefore, the possible labels were: contradiction, neutral, or entailment.
The list above is non-exhaustive. This paper aims to evaluate text-based sampling methods (see Section 3) to gather more useful texts for the fine-tuning process. Considering the above loss functions, we selected BATL, CTL, OCL, and SL. CSL and MNRL were not selected because CSL depends on a similarity measure for the text pairs and, therefore, would require extra information for this calculation. MNRL, on the other hand, requires positive pairs, or triplets, with a negative anchor. Although we could leverage the classes to generate positive pairs, choosing good-quality anchors can be challenging and require deeper analysis for an assertive selection.
@@ 3 Text-based Sampling Methods
In addition to each sampling method, except for the random sampling, we evaluate an extra scenario leveraging the text labels (classes). Therefore, the sampling methods correspond to their original version and the version that accounts for the class. Alg. 1 provides the weighted sampling pseudocode. We highlight that, optionally, the observed classes’ frequencies can be used as an argument for the WeightedSampling function. These frequencies can also be calculated directly from the buffer, using the attribute class. The algorithm runs according to the following steps: (1) the buffer containing the stored items from the stream is iterated; (2) each item has its weight calculated, depending on the chosen sampling method; (3) if the classes’ frequencies are considered, then the items’ probabilities of less frequent classes are increased proportionally; (4) the weights are normalized; and (5) $n_{s}$ instances are sampled from the buffer.
@@
  • WordpieceToken ratio sampling: we propose this sampling method based on the ratio between wordpieces and tokens. Wordpiece is a technique some BERT-based models use to handle out-of-vocabulary (OOV) tokens [7]. For example, the word institutionalization is partitioned into two wordpieces: [‘institutional’, ‘##ization’], where ## means that there is a previous partition. To the best of our knowledge, this is the first time the ratio between Wordpieces and tokens is used as a proxy for text sampling. Therefore, the rationale behind this sampling method is that the bigger the ratio between wordpieces and tokens, the bigger the number of unknown words (by the language model) in the text. That is, more wordpieces had to be used to represent words. Thus, sampling texts by weighting the WordpieceToken ratio may retrieve texts with more useful information for the fine-tuning process.
@@ 4.1.1 Datasets
An important characteristic regarding the above datasets is the presence of a timestamp field, which is crucial for text stream simulation. The data distributions are presented in Fig. 1. Noticeably, the datasets are imbalanced, reflecting the nature of real-world data.
@@
  • Classifier: we selected the Incremental Support Vector Machine (ISVM) as a classifier since several works reached relevant results using it, e.g., [6];
@@ 4.4 Evaluation Metrics
Given the class imbalance of the datasets used in experimentation, results were reported regarding Macro F1 Score and elapsed time. In particular, the Macro F1 Score averages the harmonic mean of precision and recall obtained per class.
@@ 4.5 Results
Table 1 condenses the results obtained. The bold values are the best per combination of dataset/sample size/loss function (LF), i.e., per row. The values in yellow and green are the best Macro F1 scores and elapsed times per dataset/sample size. WordPieceToken ratio (class) obtained the best Macro F1-Scores in 5 out of 8 dataset/sample size pairs, and WordPieceToken ratio in 1 out of 8. Sampling at random was the fastest in 6 out of 8 dataset/sample size pairs. Furthermore, the best Macro F1-Scores increase with the sample size.
Although elapsed time is important and variations followed similar patterns, i.e., the higher the sample size, the longer the elapsed time, the differences between them show that they may not impede fine-tuning in text stream scenarios, also depending on the hardware. The differences between maximum and minimum elapsed times are 187.07 seconds for Airbnb and 73.54 seconds for Yelp, which represent 37% and 13% of their respective average elapsed times.
@@
5 Conclusion
This paper evaluates four different sampling methods, i.e., random sampling, length sampling, TF-IDF sampling, and WordPieceToken ratio sampling (proposed in this paper), where the latter three had a version that accounted for instances’ classes, aiming at sampling important, informative texts from the buffer. Therefore, in total, seven sampling methods were evaluated. Combined with these text sampling methods, four loss functions were assessed, i.e., Batch All Triplets loss (BATL), Contrastive Tension loss (CTL), Online Contrastive loss (OCL), and Softmax loss (SL).
We observed that the chosen loss function plays a crucial role in improving the performance of the text classification task. Considering the scenario developed in this paper, CTL and OCL functions were insufficient to keep good performance levels. On the other hand, BATL and SL functions could help maintain interesting performance levels, sometimes above the SBERT without update (our baseline), suggesting the SBERT can benefit from these loss functions. In addition, the elapsed times were comparable to the baseline. Therefore, BATL and SL were the most suitable loss functions for text stream classification among those evaluated in this paper. However, one must be careful when employing this approach because fine-tuning can create a bottleneck in real-time applications, depending on the latency. Finally, text sampling methods are also crucial in fine-tuning. Our experiments suggest that the proposed WordPieceToken ratio method, especially leveraging the instances’ classes, can retrieve more informative texts and favor the machine learning model’s performance after fine-tuning. In future works, we intend to (a) extend to different stream mining tasks and (b) identify proper moments to automatically trigger the fine-tuning.
@@ Acknowledgements
Cristiano Mesquita Garcia is a grantee of a Doutorado Sanduíche scholarship provided by Fundação Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES). This paper is also a byproduct of a CNPq project coordinated by Prof. Jean Paul Barddal under grant number 409371/2022-0.

### v2

@@ Abstract
The proliferation of textual data on the Internet presents a unique opportunity for institutions and companies to monitor public opinion about their services and products. Given the rapid generation of such data, the text stream mining setting, which handles sequentially arriving, potentially infinite text streams, is often more suitable than traditional batch learning. While pre-trained language models are commonly employed for their high-quality text vectorization capabilities in streaming contexts, they face challenges adapting to concept drift—the phenomenon where the data distribution changes over time, adversely affecting model performance. Addressing the issue of concept drift, this study explores the efficacy of seven text sampling methods designed to fine-tune language models, thereby mitigating performance degradation selectively. We precisely assess the impact of these methods on fine-tuning the SBERT model using four different loss functions. Our evaluation, focused on Macro F1-score and elapsed time, employs two text stream datasets and an incremental SVM classifier to benchmark performance. Our findings indicate that Softmax loss and Batch All Triplets loss are particularly effective for text stream classification, demonstrating that larger sample sizes correlate with improved macro F1 scores. Notably, our proposed WordPieceToken ratio sampling method significantly enhances performance with the identified loss functions, surpassing baseline results.
@@ 1 Introduction
Automatically learning from textual data leveraging machine learning mechanisms brings several challenges in batch processing, such as text standardization and vectorization. Text vectorization is essential since most machine learning methods expect numeric vectors as input. Traditional vector representations, such as Bag-of-Words (BOW) [12] and Term Frequency - Inverse of Document Frequency (TF-IDF) [21], can generate very-high-dimensional vectors, which can be disadvantageous to the machine learning model, increasing the computational cost.
@@
The contributions of this paper are four-fold: (a) an extensive comparison among text sampling methods for fine-tuning purposes; (b) an analysis of the impact of the sampling methods considering the text stream setting; (c) an evaluation of loss functions for fine-tuning SBERT, and (d) a novel textual sampling method based on the ratio between Wordpieces and tokens of a text. The term Wordpieces represents a subword partition system present in BERT [8] that allows handling out-of-vocabulary tokens.
@@ 2 Background
This section presents core concepts for understanding this paper. We introduce text stream mining, SentenceBERT, and its fine-tuning process.
@@ 2.1 Text Stream Mining
According to Bifet et al. [3], data streams are “an abstraction that allows real-time analytics”. In a data stream, the items arrive individually or in small sequential batches, and the stream itself can be infinite [3, 9]. Different learning approaches have been developed to learn from data streams, including, for instance [9], the ability to learn incrementally, single-pass operations, elimination of input data as soon as possible after learning from it, and consumption of modest resources, i.e., processing power and time.
@@
Frequently, text-related approaches leverage pre-trained language models [10, 25]. Using pre-trained models can help save time since training a language model from scratch is computationally costly in time and resources [23]. In addition, the language model can easily be reused in different scenarios. However, an important drawback is that texts generally suffer from concept drift. Concept drift is a phenomenon frequently observed in real-world datasets and corresponds to changes in data distribution over time [9, 10]. Leveraging pre-trained language models without accounting for concept drift can decrease performance in the downstream task since the texts would be represented using relatively old representations [10].
@@ 2.2.1 Fine-tuning, data preparation, and loss functions
  • Softmax loss (SL) receives text pairs and a label indicating their relationship. Reimers and Gurevych [20] leveraged this loss function for natural language inference, and therefore, the possible labels were contradiction, neutral, or entailment.
The list above is non-exhaustive. This paper aims to evaluate text-based sampling methods (see Section 3) to gather more useful texts for fine-tuning. Considering the above loss functions, we selected BATL, CTL, OCL, and SL. We did not select CSL and MNRL because CSL depends on a similarity measure for the text pairs and would require extra information for this calculation. MNRL, on the other hand, requires positive pairs, or triplets, with a negative anchor. Although we could leverage the classes to generate positive pairs, choosing good-quality anchors can be challenging and require deeper analysis for an assertive selection.
@@ 3 Text-based Sampling Methods
In addition to each sampling method, except for the random sampling, we evaluate an extra scenario leveraging the text labels (classes). Therefore, the sampling methods correspond to their original version and the version that accounts for the class. Algorithm 1 provides the weighted sampling pseudocode. We highlight that, optionally, the observed classes’ frequencies can be used as an argument for the WeightedSampling function. These frequencies can also be calculated directly from the buffer, using the attribute class. The algorithm runs according to the following steps: (1) the buffer containing the stored items from the stream is iterated; (2) each item has its weight calculated, depending on the chosen sampling method; (3) if the classes’ frequencies are considered, then the items’ probabilities of less frequent classes are increased proportionally; (4) the weights are normalized; and (5) $n_{s}$ instances are sampled from the buffer.
@@
In addition to these sampling methods, we propose a novel text sampling method named WordpieceToken ratio.
3.1 WordpieceToken ratio sampling
This paper proposes a novel sampling method based on the ratio between wordpieces and tokens. Wordpiece is a technique BERT-based models use to handle out-of-vocabulary (OOV) tokens [8]. For example, the word institutionalization could be partitioned into two wordpieces: [‘institutional’, ‘##ization’], where ## means that there is a previous partition.
The rationale behind this sampling method is that the bigger the ratio between wordpieces and tokens, the bigger the number of unknown words (by the language model) in the text. The ratio is calculated per instance. For example, let the text instance be: “Extramedullary toxicity was limited to hypothyroidism”, a 6-token sentence. When applied to a BERT tokenizer, it is split into [‘extra’, ‘##med’, ‘##ulla’, ‘##ry’, ‘toxicity’, ‘was’, ‘limited’, ‘to’, ‘h’, ‘##yp’, ‘##oth’, ‘##yr’, ‘##oid’, ‘##ism’], resulting in 14 wordpieces. Thus, this instance’s weight is $\frac{14}{6}=2.33$. After all the text instances from the buffer have their weights calculated, several instances are sampled considering the instances’ weights and class frequencies according to Algorithm 1.
Therefore, we hypothesize that sampling texts by weighting the WordpieceToken ratio may retrieve texts with more useful information for the fine-tuning process.
@@ 4.1.1 Datasets
An essential characteristic regarding the above datasets is the presence of a timestamp, which is crucial for text stream simulation. The data distributions are presented in Fig. 1. Noticeably, the datasets are imbalanced, reflecting the nature of real-world data.
@@
  • Classifier: we selected the Incremental Support Vector Machine (ISVM) as the classifier because it is updated incrementally according to the arrival of new data. As its batch counterpart [6], ISVM calculates an optimal hyperplane between instances of two distinct classes, and the surrounding instances are assumed as support vectors [7].
@@ 4.4 Evaluation Metrics
Given the class imbalance of the datasets used in experimentation, results regarding Macro F1 Score and elapsed time were reported. In particular, the Macro F1 Score averages the harmonic mean of precision and recall obtained per class.
@@ 4.5 Results
Table 1 condenses the results obtained. The bold values are the best per dataset/sample size/loss function (LF) combination, i.e., per row. The values in yellow and green are the best Macro F1 scores and elapsed times per dataset/sample size. WordPieceToken ratio (class) obtained the best Macro F1-Scores in 5 out of 8 dataset/sample size pairs, and WordPieceToken ratio in 1 out of 8. Sampling at random was the fastest in 6 out of 8 dataset/sample size pairs. Furthermore, the best Macro F1-Scores increase with the sample size.
Although elapsed time is important and variations followed similar patterns, i.e., the higher the sample size, the longer the elapsed time, the differences between them show that they may not impede fine-tuning in text stream scenarios, depending on the hardware. The differences between maximum and minimum elapsed times are 187.07 seconds for Airbnb and 73.54 seconds for Yelp, which represent 37% and 13% of their respective average elapsed times.
5 Conclusion
This paper evaluates four different sampling methods, i.e., random sampling, length sampling, TF-IDF sampling, and WordPieceToken ratio sampling (proposed in this paper), where the latter three had a version that accounted for instances’ classes, aiming at sampling important, informative texts from the buffer. Four loss functions were assessed in combination with the text sampling methods, i.e., Batch All Triplets loss (BATL), Contrastive Tension loss (CTL), Online Contrastive loss (OCL), and Softmax loss (SL).
We observed that the loss function plays a crucial role in improving the performance of the text classification task. Considering the scenarios assessed, CTL and OCL functions were insufficient to achieve satisfactory performance levels. On the other hand, BATL and SL functions aided in maintaining interesting performance levels, sometimes above the SBERT without update (our baseline), suggesting the SBERT can benefit from these loss functions. In addition, the elapsed times were comparable to the baseline. Therefore, BATL and SL were the most suitable functions for text stream classification among those evaluated in this paper. However, one must be careful when employing this approach because fine-tuning can create a bottleneck in real-time applications. Finally, text sampling methods are also crucial in fine-tuning. Our experiments suggest that the proposed WordPieceToken ratio method, especially leveraging the instances’ classes, can retrieve more informative texts and favor the machine learning model’s performance after fine-tuning.
@@ 5 Conclusion
In future works, we intend to (a) extend to different stream mining tasks, (b) identify proper moments to automatically trigger the fine-tuning, (c) evaluate the balance between the sampling method and the resources consumption, and (d) evaluate other pre-trained BERT-based models with the combinations of loss functions and sampling method used in this paper.
@@ Acknowledgements
Cristiano Mesquita Garcia is a grantee of a Doutorado Sanduíche scholarship provided by Fundação Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES). This paper is also a byproduct of a CNPq project coordinated by Prof. Jean Paul Barddal under grant number 409371/2022-0 and is tied to CNPq International #441610/2023-4.